OpenStudy (anonymous):

Complete the square to form a perfect square trinomial. x² + 12x + _ a) 36 b) 16 c) 42 d) 24

OpenStudy (austinl):

I would have to say option A. \[(x+6)(x+6) = x^2+12x+36\]

OpenStudy (uri):

Yeah 36 because 6+6=12 and 6x6=36.Correct!